OpenStudy (anonymous):

Evaluate 2a - b for a = 2 and b = 3.

jimthompson5910 (jim_thompson5910):

a = 2 so you replace EVERY copy of 'a' with 2 b = 3 so you replace EVERY copy of 'b' with 3

jimthompson5910 (jim_thompson5910):

2a-b turns into 2*(2) - (3) = 4 - 3 = 1 therefore, 2a-b = 1 when a = 2 and b = 3

OpenStudy (anonymous):

thanks